# Quantum optical coherence theory based on Feynman’s path integral

### v1

@@ IV.2.2 Second-order interference of two classical light beams
The same method can be employed to calculate the second-order interference of two classical light beams in a HOM interferometer. Assume S1 and S2 are two independent and identical single-mode continuous-wave lasers. The polarizations and intensities of these two laser light beams are the same. There are three different ways to trigger a two-photon coincidence count in this case. The first one is these two photons are both emitted by S1. The probability is 1/4. There is one path and the corresponding two-photon probability amplitude is $A_{11,12}$. The second way is these two photons are both emitted by S2. The probability is 1/4. There is one path and the corresponding two-photon probability amplitude is $A_{21,22}$. The third way is the detected two photons are emitted by S1 and S2 one each. The probability is 1/2 and there are two different paths in this way. The corresponding two-photon probability amplitudes are $A_{11,22}$ and $A_{12,21}$. If these four different paths are indistinguishable, the two-photon probability function is
@@
where $\varphi_{11,12}$ is the phase related to the initial phase of these two photons and extra $\pi/2$ phase due to the reflection on BS. The last $1/\sqrt{4}$ in Eq. (74) is the product of $1/\sqrt{2}$ and $1/\sqrt{2}$, where the first $1/\sqrt{2}$ is due to the probability of these two paths equals 1/2 and the second $1/\sqrt{2}$ is the normalization parameter. Table 1 summarizes the phases for these four different paths, where $\varphi_{1}$ and $\varphi_{2}$ are the initial phases of photons emitted by S1 and S2, respectively.
@@
Assuming S1 and S2 are point light sources, temporal two-photon probability function can be obtained via simplifying Eq. (IV.2.2)
@@
where $\Delta\omega=\omega_{1}-\omega_{2}$ is frequency difference between these two laser light beams. There is second-order interference pattern by superposing two independent laser light beams. Unlike the transient first-order interference pattern of two independent laser light beams, the second-order interference pattern in Eq. (77) is observed over long time average.
Figure 20: Temporal second-order interference pattern of two independent laser light beams in a HOM interferometer Liu et al. (2016)
Figure 20 shows the temporal second-order interference pattern with two tunable single-mode continuous-wave lasers. The frequency bandwidth of laser light is 100 kHz. The frequency difference between these two lasers are 189 MHz, 461 MHz, 1.11 GHz, and 3.0 GHz in Figs. 20(a)-(d), respectively. Periodic interference pattern is observed in Figs. 20(a)-(b). There are also periodic structure in Figs. 20(c)-(d). The reason why it can not be observed is the period of the interference pattern is too small to observe in the present time scale Liu et al. (2016).
The same method can be employed to calculate the second-order interference of two independent thermal and laser light beams. Assume S1 is single-mode continuous-wave laser and S2 is thermal light source in Fig. 18. The intensities, polarizations, frequencies of the light beams emitted by these two sources are the same. There are three different ways for D1 and D2 detecting a two-photon coincidence count. The first way is these two photons are emitted by laser S1. The probability is 1/4 and there is one path, $A_{11,12}$. The second way is these two photons are emitted by thermal light source S2. The probability is 1/4 and there are two paths, $A_{2a1,2b2}$ and $A_{2a2,2b1}$. The third way is these two photons are emitted by S1 and S2 each. The probability is 1/2 and there are two different paths, $A_{11,22}$ and $A_{12,21}$. If the five different paths are indistinguishable, the two-photon probability function equals
@@
Figure 21: Second-order interference pattern of thermal and laser light (b) and the second-order coherence function of laser and thermal light beams in the interferometer alone (a) Liu et al. (2014b).
Figure 21(b) shows the second-order interference pattern of thermal and laser light beams in a HOM interferometer, which is consistent with Eq. (IV.2.2). For comparison, Fig. 21(a) shows the measured second-order coherence functions when one of the incidental light beams is blocked, which is equivalent to laser and thermal light in a HBT interferometer Liu et al. (2014b).
The second-order interference between two independent thermal light beams Liu et al. (2013), classical and nonclassical light beams Hong et al. (1988) and other two independent light beams can also be calculated in the same method above. The main difference between the second-order of two independent classical and nonclassical light beams is that the visibility of the second-order interference pattern of two independent nonclassicla light beams can reach 100%, while the maximal visibility of the second-order interference pattern between two independent classical light beam is 50% Mandel (1983). The reason is clearly shown in Eqs. (70), (74), and (78). In the second-order interference of two independent classical light beams, it is impossible to distinguish one source emitting two photons from two sources emitting a photon each, which will contribute to the background coincidence counts. If special two-photon coincidence count system can distinguish these two types of coincidence count, the visibility of second-order interference pattern of two independent classical light beams can exceed 50%. For instance, Kaltenbaek et al. observed the visibility of second-order interference pattern with two independent laser light beams exceeding 85% by employing sum frequency generation to distinguish these two different types of coincidence count. With two non-independent classical light beams and normal two-photon coincidence count system, Sadana et al. observed near 100% visibility second-order interference pattern Sadana et al. (2019). The results do not conflict with Mandel’s conclusion due to Mandel’s conclusion is valid for two independent light beams and normal two-photon coincidence count detection system Mandel (1983).
@@ IV.3 Second-order interference of multiple light beams
Feyman’s path integral can be employed to calculate the second-order interference of more than two independent light beams. Figure 22 shows the second-order interference of three independent light beams. S1, S2 and S3 are three independent light sources. D1 and D2 are two single-photon detectors. The output of two detectors are sent into a two-photon coincidence count system to measure the second-order correlation, which is not shown in the figure.
Figure 22: Second-order interference of three independent light beams
Assume S1, S2 and S3 are three independent and identical single-photon sources. There are three different ways to trigger a two-photon coincidence count. The first one is the two detected photons are emitted by S1 and S2. The probability is 1/3. There are two different paths and the corresponding two-photon probability amplitudes are $A_{11,22}$ and $A_{12,21}$. The second way is these two photons are emitted by S1 and S3. The probability is 1/3. There are two paths and the corresponding two-photon probability amplitudes are $A_{11,32}$ and $A_{12,31}$. The third way is these two photons are emitted by S2 and S3. The probability is 1/3. There are two paths and the corresponding two-photon probability amplitudes are $A_{21,32}$ and $A_{22,31}$. If only one single-photon source emit a photon, it is impossible to trigger a two-photon coincidence count. If two single-photon sources emit photons simultaneously, these three different ways are distinguishable. One can measure the status of the single-photon source to ensure which two sources emitted photons. The probability for three independent single-photon sources simultaneously emitting photons is much smaller than the one of two single-photon sources simultaneously emitting photons, which will be ignored in the calculation.
@@
With the same method above, Eq. (81) can be simplified by considering the three single-photon sources are independent,
@@
$=$ $\langle|K_{11}K_{22}+K_{12}K_{21}|^{2}\rangle+\langle|K_{11}K_{32}+K_{12}K_{31}|^{2}\rangle$
@@
where $K_{11}$ is the photon’s Feynman propagator from S1 to D1 and the meanings of other symbols are defined similarly. Assuming S1, S2 and S3 are point source and in the same source plane, the distance between the source and the observation planes is much larger than one between sources, the spatial two-photon probability function can be obtained via Eq. (82)
@@
where paraxial approximation and one-dimension have been assumed, $\lambda$ is the wavelength of light and single-frequency is employed to simplify the calculation, $d_{12}$ is the distance between S1 and S2 and relative positions of these three sources are shown in Fig. 8. The meanings of other symbols are defined similarly.
Figure 23: Second-order interference pattern of three independent light beams. $x_{1}-x_{2}$ is the relative position of D1 and D2.
Figure 23 shows the second-order interference pattern of three single-photon sources shown in Fig. 22. The patterns will change as the distances between these three sources change. Compared to the first-order interference pattern of three light beams shown in Fig. 9, the visibility of the second-order interference pattern with three independent light beams can not reach 100% even when single-photon states are employed.
The second-order interference of three independent laser light beams can also be calculated in the same method. Assume S1, S2 and S3 are three independent and single-mode continuous-wave lasers and the polarizations of these three laser light beams are identical. There are six different ways to trigger a two-photon coincidence count in Fig. 22. The first way is these two photons are both emitted by S1. The probability is $I_{1}^{2}/(I_{1}+I_{2}+I_{3})^{2}$, where $I_{1}$, $I_{2}$, and $I_{3}$ are the intensities of laser light beams emitted by S1, S2 and S3, respectively. There is one path and the corresponding two-photon probability amplitude is $A_{11,12}$. The second ways is these two photons are both emitted by S2. The probability is $I_{2}^{2}/(I_{1}+I_{2}+I_{3})^{2}$. There is one path and the two-photon probability amplitude is $A_{21,22}$. The third way is these two photons are both emitted by S3. The probability is $I_{3}^{2}/(I_{1}+I_{2}+I_{3})^{2}$. There is one path and the probability amplitude is $A_{31,32}$. The fourth way is these two photons are emitted by S1 and S2 one each. The probability is $2I_{1}I_{2}/(I_{1}+I_{2}+I_{3})^{2}$. There are two paths and the probability amplitudes are $A_{11,22}$ and $A_{12,21}$. The fifth way is these two photons are emitted by S1 and S3 one each. The probability is $2I_{1}I_{3}/(I_{1}+I_{2}+I_{3})^{2}$ and the two paths are $A_{11,32}$ and $A_{12,31}$. The sixth way is these two photons are emitted by S2 and S3 one each. The probability is $2I_{2}I_{3}/(I_{1}+I_{2}+I_{3})^{2}$ and the two paths are $A_{21,32}$ and $A_{22,31}$. If all the nine different paths to trigger a two-photon coincidence count are indistinguishable, the two-photon probability function is
@@
where the common factor $1/(I_{1}+I_{2}+I_{3})$ has been ignored, $\sqrt{I_{1}I_{2}}=\sqrt{2I_{1}I_{2}}\times 1/\sqrt{2}$.
@@
$+I_{1}I_{3}\langle|K_{11}K_{32}+K_{12}K_{31}|^{2}\rangle+I_{2}I_{3}\langle|K_{21}K_{32}+K_{22}K_{31}|^{2}\rangle.$
Assuming the intensities of these three light beams are equal and single frequency light is emitted, Eq. (IV.3) can be simplified as
$P^{(2)}(x_{1}-x_{2})$
@@
where paraxial approximation and one-dimension are assumed to simplify the calculations. Comparing the second-order interference pattern of three independent single-photon states, Eq. (IV.3) and the second-order interference pattern of three independent single-mode continuous-wave lasers, the interference patterns are the same except the constant changes from 3 to 4.5. The reason for this difference is that laser light can not avoid two photons emitted from the same source with normal two-photon coincidence count detection system, which is exactly the same as the one for the difference between the visibility of the second-order interference pattern of two independent classical and non-classical light beams Mandel (1983).
@@ IV.4 Subwavelength interference
In the above parts of Sect. IV, we have introduced how to employ Feynman’s path integral to calculate the second-order interference of one, two, and multiple light beams. As an example of how quantum optical coherence theory based on Feynman’s path integral will help us to understand the physics of optical coherence, subwavelength interference is analyzed in Feynman’s path integral.
Taking the first- and second-order interference of light emitted by two independent sources shown in Fig. 5 as an example to illustrate subwavelength interference. The transient first-order interference pattern of two independent single-mode continuous-wave lasers is
@@
Putting two detectors in the detection plane in Fig. 5, the second-order interference pattern of two independent single-mode continuous-wave lasers is proportional to
@@
Comparing Eq. (88) with Eq. (87) , if these two detectors are moving in the opposite direction with the same speed, the period of the second-order interference pattern will be half of the one of the first-order interference pattern, which is called subwavelength interference.
@@
Subwavelength interference can be understood better with the concept of de Broglie wavelength. The de Broglie wavelength of a photon equals $h/p$, where $h$ is Planck constant and $p$ is the momentum of the photon. If two photons can be treated as a whole, the corresponding de Broglie wavelength equals $h/(2p)$, which is only half of the wavelength of one photon. The same goes for the de Broglie wavelength of $N$ photon equals $h/(Np)$, where $N$ is a positive integer larger than 1 Jacobson et al. (1995). Two-photon subwavelength interference was first observed with entangled photon pairs Fonseca et al. (1999); Boto et al. (2000); D’Angelo et al. (2001). Then it was found that subwavelength interference can be observed with thermal light Scarcelli et al. (2004); Xiong et al. (2005); Zhai et al. (2005) and laser light Liu and Zhang (2010).
Figure 24: Scheme for subwavelength interference with two laser light beams Liu and Zhang (2010)
@@
Figure 24 shows the scheme for observing two-photon subwavelength with two independent laser light beams Liu and Zhang (2010), which is equivalent to a two point source interferometer. Two independent laser light beams was obtained by splitting a single-mode continuous-wave laser light beam by a beam splitter (BS1) and randomize the phase of one beam by a mirror mounted on a piezoelectric ceramic (PM). One beam splitter (BS2), two single-photon detectors (D1 and D2), and a two-photon coincidence count detection system (CC) are employed to measure the second-order interference pattern.
Figure 25: The observed normal second-order interference pattern Liu and Zhang (2010).
@@
Figure 25(b) shows the normal second-order interference pattern by fixing the position of D2 and scanning the position of D1. The period of the observed second-order interference pattern is 800 $\mu$m. The visibility is $(46\pm 2)\%$, which does not exceed 50%. Figure 25(a) shows the single-photon counting rates observed by two single-photon detectors, respectively. There is no first-order interference pattern observed.
@@
Figure 26: Two-photon subwavelength interference pattern Liu and Zhang (2010).
@@
Figure 26(b) shows the observed two-photon subwavelength interference pattern by by scanning the positions of D1 and D2 in the opposite directions. The period of the second-order interference pattern is 400 $\mu$m, which is exactly half of the one shown in Fig. 25(b) and indicates subwavelength interference is observed. The visibility of the second-order interference pattern equals $(48\pm 4)\%$. One may argue that it is meaningless to observe subwavelength by scanning two detectors in the opposite directions Liu et al. (2014c). However, one can obtain subwavelength interference with two detectors scanning in the same direction for classical light with suitable changes on the experimental scheme Cao et al. (2010), which is similar as the one with entangled photon pairs Fonseca et al. (1999); D’Angelo et al. (2001). As shown below, there is no such big difference between the observed subwavelength interference patterns for detectors scanning in the same direction and opposite directions.
@@
In order to see the physics of subwavelength interference, take the second-order interference of two single-mode continuous-wave laser with the same phase as an example. Assume the intensities and polarizations of these two laser light beams are the same. There are four different paths to trigger a two-photon coincidence count in Fig. 24. If these four difference paths are indistinguishable , the two-photon probability function is
@@
$=$ $\langle|A_{11,12}+A_{21,22}+A_{11,22}+A_{12,21}|^{2}\rangle$
$=$ $|e^{i2\varphi}(K_{11}K_{12}+K_{21}K_{22}+K_{11}K_{22}+K_{12}K_{21}|^{2}\rangle$
@@
where $I_{1}=|K_{11}|^{2}=|K_{21}|^{2}$ and $I_{2}=|K_{12}|^{2}=|K_{22}|^{2}$, which equals half of the intensities detected by D1 and D2, respectively. Assume point light sources and single-frequency light beams are employed, Eq. (90) can be simplified as Liu and Zhang (2010)
@@
The last two lines on the right hand side of Eq. (91) corresponds to two-photon subwavelength when two detectors scanning in the same direction and opposition directions, respectively. $2I_{1}I_{2}\cos\frac{2\pi d}{\lambda L}(x_{1}+x_{2})$ corresponds to $K_{11}K_{12}K_{21}^{*}K_{22}^{*}+K_{21}K_{22}K_{11}^{*}K_{12}^{*}$ in Eq. (90), which is a result of two-photon interference for these two photons emitted by the same source. It is consistent with the condition for observing subwavelength interference with entangled photon pairs Fonseca et al. (1999); D’Angelo et al. (2001). $2I_{1}I_{2}\cos\frac{2\pi d}{\lambda L}(x_{1}-x_{2})$ corresponds to $K_{11}K_{22}K_{12}^{*}K_{21}^{*}+K_{12}K_{21}K_{11}^{*}K_{22}^{*}$ in Eq. (90), which is a result of two-photon interference for these two photons emitted by one source each. If one can eliminate other terms and only keep one of the last two terms in Eq. (91), one can observe subwavelength interference. The experiment reported in Ref. Liu and Zhang (2010) employed random phases between these two lasers to eliminate all the other terms except the constant and the last term in Eq. (91). Subwavelength interference is observed when two detectors are scanning in the opposite directions. Equations (90) and (91) shows that there is no intrinsic difference between the subwavelength interference when these two detectors scanning in the same direction and opposite directions. They are results of two-photon interference with different two-photon paths.

### v2

@@ IV.2.2 Second-order interference of two classical light beams
The same method can be employed to calculate the second-order interference of two classical light beams in a HOM interferometer. Assume S1 and S2 are two independent and identical single-mode continuous-wave lasers. The polarizations and intensities of these two laser light beams are the same. There are three different ways to trigger a two-photon coincidence count in this case. The first way is these two photons are both emitted by S1. The probability is 1/4. There is one path and the corresponding two-photon probability amplitude is $A_{11,12}$. The second way is these two photons are both emitted by S2. The probability is 1/4. There is one path and the corresponding two-photon probability amplitude is $A_{21,22}$. The third way is the detected two photons are emitted by S1 and S2 one each. The probability is 1/2 and there are two different paths. The corresponding two-photon probability amplitudes are $A_{11,22}$ and $A_{12,21}$. If these four different paths to trigger a two-photon coincidence count are indistinguishable, the two-photon probability function is
@@
where $\varphi_{11,12}$ is the phase related to the initial phase of these two photons and extra $\pi/2$ phase due to the reflection on BS. The last $1/\sqrt{4}$ in Eq. (74) is the product of $1/\sqrt{2}$ times $1/\sqrt{2}$, where the first $1/\sqrt{2}$ is due to the probability of these two paths equals 1/2 and the second $1/\sqrt{2}$ is the normalization parameter. Table 1 summarizes the phases for these four different paths, where $\varphi_{1}$ and $\varphi_{2}$ are the initial phases of photons emitted by S1 and S2, respectively.
@@
Assuming S1 and S2 are point laser light sources, temporal two-photon probability function can be obtained by simplifying Eq. (IV.2.2),
@@
where $\Delta\omega=\omega_{1}-\omega_{2}$ is frequency difference between these two laser light beams. There is second-order interference pattern by superposing two independent laser light beams. Unlike the transient first-order interference pattern of two independent laser light beams is observed within the coherence time, the second-order interference pattern in Eq. (77) is observed over long time average.
Figure 20: Temporal second-order interference patterns of two independent laser light beams in a HOM interferometer Liu et al. (2016)
Figure 20 shows the temporal second-order interference patterns with two tunable single-mode continuous-wave lasers. The frequency bandwidth of laser light is 100 kHz. The frequency difference between these two lasers are 189 MHz, 461 MHz, 1.11 GHz, and 3.0 GHz in Figs. 20(a)-(d), respectively. Periodic interference patterns are observed in Figs. 20(a)-(b). There are also periodic structure in Figs. 20(c)-(d). The reason why it can not be observed is that the period of the interference pattern is too small to be observed in the time scale Liu et al. (2016).
The same method can be employed to calculate the second-order interference of two independent thermal and laser light beams. Assume S1 is a single-mode continuous-wave laser and S2 is a thermal light source in Fig. 18. The intensities, polarizations, and frequencies of the light beams emitted by these two sources are the same. There are three different ways for D1 and D2 detecting a two-photon coincidence count. The first way is these two photons are emitted by laser, S1. The probability is 1/4 and there is one path, $A_{11,12}$. The second way is these two photons are emitted by thermal light source, S2. The probability is 1/4 and there are two paths, $A_{2a1,2b2}$ and $A_{2a2,2b1}$. The third way is these two photons are emitted by S1 and S2 each. The probability is 1/2 and there are two different paths, $A_{11,22}$ and $A_{12,21}$. If these five different paths are indistinguishable, the two-photon probability function equals
@@
Figure 21: Second-order interference pattern of thermal and laser light (b) and the second-order coherence functions of laser and thermal light beams in the interferometer alone (a) Liu et al. (2014b).
Figure 21(b) shows the spatial second-order interference pattern of thermal and laser light beams in a HOM interferometer, which is consistent with Eq. (IV.2.2). For comparison, Fig. 21(a) shows the measured second-order coherence functions when one of the incidental light beams is blocked, which is equivalent to laser and thermal light in a HBT interferometer Liu et al. (2014b).
The second-order interference of two independent thermal light beams Liu et al. (2013), classical and non-classical light beams Hong et al. (1988) and other two independent light beams can also be calculated in the same method above. The main difference between the second-order of two independent classical and non-classical light beams is that the visibility of the second-order interference pattern of two independent non-classical light beams can reach 100%, while the maximal visibility of the second-order interference pattern of two independent classical light beam is 50% Mandel (1983). The reason is clearly illustrated by Eqs. (70), (74), and (78). In the second-order interference of two independent classical light beams, it is impossible to distinguish two sources emitting a photon each from one source emitting two photons, which will contribute to the background coincidence counts. If special two-photon coincidence count detection system can distinguish these two types of coincidence count, the visibility of the second-order interference pattern of two independent classical light beams can exceed 50%. For instance, Kaltenbaek et al. observed the visibility of second-order interference pattern with two independent laser light beams exceeding 85% by employing sum frequency generation to distinguish these two different types of coincidence count Kaltenbaek et al. (2008). With two non-independent classical light beams and normal two-photon coincidence count system, Sadana et al. observed the second-order interference pattern with near 100% visibility Sadana et al. (2019). Their results Sadana et al. (2019); Kaltenbaek et al. (2008) do not conflict with Mandel’s conclusion due to Mandel’s conclusion is valid for two independent light beams and normal two-photon coincidence count detection system Mandel (1983).
@@ IV.3 Second-order interference of multiple light beams
Feyman’s path integral can be employed to calculate the second-order interference of more than two independent light beams. Figure 22 shows the scheme for the second-order interference of three independent light beams. S1, S2 and S3 are three independent light sources. D1 and D2 are two single-photon detectors. The output of these two detectors are sent into a two-photon coincidence count detection system to measure the second-order correlation, which is not shown in the figure.
Figure 22: Second-order interference of three independent light beams. S1, S2 and S3 are three light sources. D1 and D2 are two single-photon detectors. The output of two detectors are sent into a two-photon coincidence count system, which is not shown.
Assume S1, S2 and S3 are three independent and identical single-photon sources. There are three different ways to trigger a two-photon coincidence count. The first way is that the two detected photons are emitted by S1 and S2. The probability is 1/3. There are two different paths and the corresponding two-photon probability amplitudes are $A_{11,22}$ and $A_{12,21}$. The second way is that these two photons are emitted by S1 and S3. The probability is 1/3. There are two paths and the corresponding two-photon probability amplitudes are $A_{11,32}$ and $A_{12,31}$. The third way is that these two photons are emitted by S2 and S3. The probability is 1/3. There are two paths and the corresponding two-photon probability amplitudes are $A_{21,32}$ and $A_{22,31}$. If only one single-photon source emits a photon, it is impossible to trigger a two-photon coincidence count. If two single-photon sources emit a photon simultaneously, these three different ways are distinguishable. One can measure the status of the single-photon sources to ensure which two sources emitted photons. The probability for three independent single-photon sources simultaneously emitting a photon is much smaller than the probability of two single-photon sources simultaneously emitting a photon. Hence three single-photon sources simultaneously emitting a photon can be ignored in the calculation.
@@
With the same method above, Eq. (81) can be simplified as,
@@
$\propto$ $\langle|K_{11}K_{22}+K_{12}K_{21}|^{2}\rangle+\langle|K_{11}K_{32}+K_{12}K_{31}|^{2}\rangle$
@@
where $K_{11}$ is the photon’s Feynman propagator from S1 to D1 and the meanings of other symbols are defined similarly. Assuming S1, S2 and S3 are point single-photon sources in the same plane, the distance between the source and the observation planes is much larger than one between sources, the spatial two-photon probability function can be obtained via Eq. (82),
@@
where paraxial approximation and one-dimension case have been assumed, $\lambda$ is the wavelength of light and single-frequency light is employed to simplify the calculation, $d_{12}$ is the distance between S1 and S2 and the relative positions of these three sources are shown in Fig. 8. The meanings of other symbols are defined similarly.
Figure 23: Second-order interference patterns of three independent light beams. $x_{1}-x_{2}$ is the transverse position difference of D1 and D2.
Figure 23 shows the second-order interference patterns of three single-photon sources. The patterns will change as the distances between these three sources change. Compared to the first-order interference pattern of three light beams shown in Fig. 9, the visibility of the second-order interference pattern with three independent light beams can not reach 100% even when non-classical light are employed.
The second-order interference of three independent laser light beams can be calculated in the same method. Assume S1, S2 and S3 are three independent and single-mode continuous-wave lasers and the polarizations of these three laser light beams are identical. There are six different ways to trigger a two-photon coincidence count in Fig. 22. The first way is that these two photons are both emitted by S1. The probability is $I_{1}^{2}/(I_{1}+I_{2}+I_{3})^{2}$, where $I_{1}$, $I_{2}$, and $I_{3}$ are the intensities of laser light beams emitted by S1, S2 and S3, respectively. There is one path and the corresponding two-photon probability amplitude is $A_{11,12}$. The second ways is that these two photons are both emitted by S2. The probability is $I_{2}^{2}/(I_{1}+I_{2}+I_{3})^{2}$. There is one path and the two-photon probability amplitude is $A_{21,22}$. The third way is that these two photons are both emitted by S3. The probability is $I_{3}^{2}/(I_{1}+I_{2}+I_{3})^{2}$. There is one path and the probability amplitude is $A_{31,32}$. The fourth way is that these two photons are emitted by S1 and S2 one each. The probability is $2I_{1}I_{2}/(I_{1}+I_{2}+I_{3})^{2}$. There are two paths and the probability amplitudes are $A_{11,22}$ and $A_{12,21}$. The fifth way is that these two photons are emitted by S1 and S3 one each. The probability is $2I_{1}I_{3}/(I_{1}+I_{2}+I_{3})^{2}$ and the two paths are $A_{11,32}$ and $A_{12,31}$. The sixth way is that these two photons are emitted by S2 and S3 one each. The probability is $2I_{2}I_{3}/(I_{1}+I_{2}+I_{3})^{2}$ and the two paths are $A_{21,32}$ and $A_{22,31}$. If all the nine different paths to trigger a two-photon coincidence count are indistinguishable, the two-photon probability function is
@@
where the common factor $1/(I_{1}+I_{2}+I_{3})$ has been ignored, $\sqrt{I_{1}I_{2}}$ equals $\sqrt{2I_{1}I_{2}}\times 1/\sqrt{2}$.
@@
$+I_{1}I_{3}\langle|K_{11}K_{32}+K_{12}K_{31}|^{2}\rangle$
$+I_{2}I_{3}\langle|K_{21}K_{32}+K_{22}K_{31}|^{2}\rangle.$
Assuming the intensities of these three light beams are equal and single-frequency light is emitted, Eq. (85) can be simplified as
$P^{(2)}(x_{1}-x_{2})$ (86)
@@
$+\cos\frac{2\pi d_{13}}{L\lambda}(x_{1}-x_{2})$
$+\cos\frac{2\pi d_{23}}{L\lambda}(x_{1}-x_{2}+\frac{d_{12}}{2}),$
where paraxial approximation and one-dimension case are assumed to simplify the calculations. The second-order interference patterns of three independent single-photon states and three independent single-mode continuous-wave lasers are the same except the constant changes from 3 to 4.5. The reason for this difference is that laser light can not avoid two photons emitted by the same source with normal two-photon coincidence count detection system, which is the same as the one for the difference between the visibilities of the second-order interference patterns of two independent classical and non-classical light beams Mandel (1983).
@@ IV.4 Subwavelength interference
In the above parts of Sect. IV, we have introduced how to employ Feynman’s path integral to calculate the second-order interference of one, two, and multiple light beams. As an example of how quantum optical coherence theory based on Feynman’s path integral can be helpful to understand the physics of optical coherence, subwavelength interference of light is analyzed in Feynman’s path integral.
Take the first- and second-order interference of light emitted by two independent sources shown in Fig. 5 as an example to illustrate subwavelength interference. The transient first-order interference pattern of two independent single-mode continuous-wave lasers is
@@
Putting two detectors in the detection plane in Fig. 5, the second-order interference pattern of two independent single-mode continuous-wave lasers is
@@
Comparing Eq. (88) with Eq. (87) , if these two detectors are moving in the opposite directions with the same speed, the period of the second-order interference pattern will be half of the period of the first-order interference pattern, which is called subwavelength interference.
@@
Subwavelength interference can be understood with the concept of de Broglie wavelength. The de Broglie wavelength of a photon equals $h/p$, where $h$ is Planck constant and $p$ is the momentum of the photon. If two photons can be treated as a whole, the corresponding de Broglie wavelength equals $h/(2p)$, which is only half of the wavelength of one photon. The de Broglie wavelength of $N$ photon equals $h/(Np)$, where $N$ is a positive integer larger than 1 Jacobson et al. (1995). Two-photon subwavelength interference was first observed with entangled photon pairs Fonseca et al. (1999); Boto et al. (2000); D’Angelo et al. (2001). Then it was found that two-photon subwavelength interference can also be observed with thermal light Scarcelli et al. (2004); Xiong et al. (2005); Zhai et al. (2005) and laser light Liu and Zhang (2010).
Figure 24: Scheme for two-photon subwavelength interference with two laser light beams Liu and Zhang (2010). Laser is a single-mode continuous-wave laser. M is a mirror. PM is mirror mounted on a piezoelectric ceramic. SA and SB are two virtual sources of source S of M and PM, respectively. The meanings of other symbols are similar as the ones above.
@@
Figure 24 shows the scheme for observing two-photon subwavelength interference with two independent laser light beams Liu and Zhang (2010), which is equivalent to a two point-source interferometer. Two independent laser light beams was obtained by splitting a single-mode continuous-wave laser light beam with a beam splitter (BS1) and randomizing the phase of one beam by a mirror mounted on a piezoelectric ceramic (PM). One beam splitter (BS2), two single-photon detectors (D1 and D2), and a two-photon coincidence count detection system (CC) are employed to measure the second-order interference pattern.
@@
Figure 25: The observed normal second-order interference pattern (b) when the position of D1 is scanning why the position of D2 is fixed Liu and Zhang (2010). R1 and R2 in (a) are single-photon counting rates of D1 and D2, respectively.
Figure 25(b) shows the normal second-order interference pattern by fixing the position of D2 and scanning the position of D1. The period of the observed second-order interference pattern is 800 $\mu$m. The visibility is $(46\pm 2)\%$, which does not exceed 50%. Figure 25(a) shows the single-photon counting rates observed by two single-photon detectors, respectively. There is no first-order interference pattern.
@@
Figure 26: Two-photon subwavelength interference pattern (b) when the positions of D1 and D2 are scanned in the opposite directions Liu and Zhang (2010). R1 and R2 in (a) are single-photon counting rates of D1 and D2, respectively.
@@
Figure 26(b) shows the observed two-photon subwavelength interference pattern by scanning D1 and D2 in the opposite directions. The period of the second-order interference pattern is 400 $\mu$m, which is exactly half of the one shown in Fig. 25(b) and indicates two-photon subwavelength interference is observed. The visibility of the second-order interference pattern equals $(48\pm 4)\%$. One may argue that it is meaningless to observe subwavelength interference by scanning two detectors in the opposite directions Liu et al. (2014c). However, two-photon subwavelength interference with two detectors scanning in the same direction for classical light can be observed with suitable changes on the experimental scheme Cao et al. (2010), which is similar as the one with entangled photon pairs Fonseca et al. (1999); D’Angelo et al. (2001). As shown below, there is no intrinsic difference between the observed subwavelength interference patterns for detectors scanning in the same direction and opposite directions.
@@
In order to understand the physics of subwavelength interference, take the second-order interference of two single-mode continuous-wave laser light beams with the same phase as an example. Assume the intensities, frequency spectra and polarizations of these two laser light beams are the same. There are four different paths to trigger a two-photon coincidence count in Fig. 24. If these four difference paths are indistinguishable , the two-photon probability function is
@@
$=$ $\langle|\frac{1}{\sqrt{4}}[A_{11,12}+A_{21,22}+A_{11,22}+A_{12,21}]|^{2}\rangle$
$\propto$ $|e^{i2\varphi}(K_{11}K_{12}+K_{21}K_{22}+K_{11}K_{22}+K_{12}K_{21}|^{2}\rangle$
@@
where $I_{1}=|K_{11}|^{2}=|K_{21}|^{2}$ and $I_{2}=|K_{12}|^{2}=|K_{22}|^{2}$, which equals half of the intensities detected by D1 and D2, respectively. Assume point laser light sources are employed, Eq. (90) can be simplified as Liu and Zhang (2010)
@@
The last two lines on the right hand side of Eq. (91) correspond to two-photon subwavelength when two detectors scanning in the same direction and opposition directions, respectively. $2I_{1}I_{2}\cos\frac{2\pi d}{\lambda L}(x_{1}+x_{2})$ corresponds to $K_{11}K_{12}K_{21}^{*}K_{22}^{*}+K_{21}K_{22}K_{11}^{*}K_{12}^{*}$ in Eq. (90), which is a result of two-photon interference for these two photons emitted by the same source. It is consistent with the condition for observing subwavelength interference with entangled photon pairs Fonseca et al. (1999); D’Angelo et al. (2001). $2I_{1}I_{2}\cos\frac{2\pi d}{\lambda L}(x_{1}-x_{2})$ corresponds to $K_{11}K_{22}K_{12}^{*}K_{21}^{*}+K_{12}K_{21}K_{11}^{*}K_{22}^{*}$ in Eq. (90), which is a result of two-photon interference for these two photons emitted by one source each. If one can eliminate all the other terms and keep only one of the last two terms in Eq. (91), one can observe subwavelength interference. The experiment reported in Ref. Liu and Zhang (2010) employed random phases between these two lasers to eliminate all the other terms except the constant and the last term in Eq. (91). Subwavelength interference is observed when these two detectors are scanning in the opposite directions. Equations (90) and (91) shows that there is no intrinsic difference between the subwavelength interference when these two detectors are scanning in the same direction and opposite directions. They are results of two-photon interference with different two-photon paths.